this is Quadratic or not Quadratic?
1/2 (h-6)=0

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

Hint: there's no 2nd degree term.

Answer by Theo(13342)

not quadratic.

here's a reference on quadratic functions.

https://www.mathsisfun.com/algebra/quadratic-equation.html

basically, it's an expression where the highest exponent is 2, and the variable with the exponent of 2 has to be in the expression.

your expression is 1/2 (h-6)=0

the highest exponent in your expression is h, which is the same as h^1 which means h raised to the exponent of 1.

that would not be quadratic.

an expression that included h^2 would be quadratic.

sometime the expression might be in the of (h - 6) * h.

while that doesn't, at first glance, look to be quadratic, when you multiply out the factors you will see that it is quadratic because it becomes h^2 - 6h.

the equation of 1/2 (h-6) = 0 becomes 1/2 * h - 3 = 0 when you multiply out the factors.

the highest exponent is 1 which is not quadratic.

to be quadratic, the highest exponent in the expression has to be 2.